# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ Appendix H Heuristic estimate of the diffusion coefficient
$\left\langle\mathrm{e}^{\mathrm{i}k\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}\eta(t-t^{\prime})}\right\rangle=\mathrm{e}^{-k^{2}D_{\theta\theta}\tau},$ (74)
@@
$\frac{\partial F}{\partial t}{}$ $=4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (75)
×$\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\mathrm{e}^{-\tau/T_{k}}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$
where we introduced the (scale-dependent) decorrelation time, ${T_{k}\!=\!1/(2D_{\theta\theta}k^{2})}$. Integrating over $\tau$, we obtain
$\frac{\partial F}{\partial t}=4\pi^{2}\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (76)
×$\times\delta_{T_{k}}\big{(}k\,\Delta\Omega[J,J_{1}]\big{)}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$
with ${\delta_{T_{k}}(\omega)}$ the broadened Dirac delta from Eq. (24). [^10] In the main text, we assumed ${T_{k}\!=\!T_{\mathrm{reg}}/k}$. This leads to the identity ${\delta_{T_{k}}(k\,\omega)\!=\!\delta_{T_{\mathrm{reg}}}(\omega)/|k|}$, similar to the one satisfied by the usual Dirac delta. From this identity results the broadened Landau equation (Eq. 25). Here, the function ${\delta_{T_{k}}(k\,\omega)}$ has a different dependence with $k$, so the broadened Landau equation (Eq. 76) is slightly different. See Chavanis and Lemou (2007) for another regularization of the Dirac delta. Explicitly, the kinetic equation reads
$\frac{\partial F}{\partial t}=4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (77)
×$\times\frac{T_{k}}{1\!+\!(k\,\Delta\Omega[J,J_{1}]\,T_{k})^{2}}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]}.$
If we let ${T_{k}\!\rightarrow\!+\infty}$, or if ${k\,\Delta\Omega\,T_{k}\!\gg\!1}$, which both amount to neglecting the fluctuations in particles individual trajectories, Eq. (77) becomes the Landau equation (Eq. 14).
However, when $J$ is close to the extremum of ${\Omega(J)}$, which we abusively take to be in $J_{0}$, we are faced with the opposite limit, namely ${k\,\Delta\Omega\,T_{k}\!\ll\!1}$. In that case, the kinetic equation (Eq. 77) reduces to
$\frac{\partial F}{\partial t}={}$ $4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
×$\times T_{k}\,\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]}.$ (78)
@@
$D=4\pi^{2}\gamma\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
×$\times\,\delta_{T_{k}}\big{(}k\,\Delta\Omega[J,J_{1}]\big{)}F(J_{1},t).$ (79)
@@
As expected, this equation reduces to the usual Landau diffusion coefficient when ${T_{k}\!\rightarrow\!+\infty}$, and scales like ${1/N}$ (via the prefactor $\gamma$). However, if we assume ${k\,\Delta\Omega\,T_{k}\!\ll\!1}$ and heuristically identify the diffusion coefficient in angles, $D_{\theta\theta}$, with the diffusion coefficient from Eq. (79), i.e. if we write ${D_{\theta\theta}\!\simeq\!D/J_{0}^{2}}$, we obtain for ${J\!\rightarrow\!J_{0}}$
$D^{2}=2\pi\gamma\,J_{0}^{2}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}|U_{k}(J,J_{1})|^{2}F(J_{1},t).$ (80)
@@
Since ${\gamma\!\sim\!1/N}$, we recover the Taylor–McNamara scaling Taylor and McNamara (1971), ${D\!\sim\!1/\sqrt{N}}$. If we account for an active fraction of particles, as in the main text, this scaling becomes ${D\!\sim\!q/\sqrt{N}}$. We note that the diffusion coefficient from Eq. (80) could be obtained numerically, and subsequently compared with measurements in numerical simulations.
@@
Along the same line, if we make the local approximation, ${F(J_{1},t)\!\simeq\!F(J_{0},t)}$, the integral in Eq. (80) can be estimated analytically when using the (unsoftened) interaction potential from Eq. (43). Overall, Eq. (80) leads to the rough estimate
$D^{2}\sim\frac{J_{0}^{3}}{8\pi}\,\gamma\,F(J_{0},t),$ (81)
@@
up to logarithmic corrections. To make a closer connection with the results of Taylor–McNamara Taylor and McNamara (1971), we introduce the original diffusion coefficient in radius $D_{rr}$, related to $D$ by ${D\!=\!D_{rr}r^{2}\!=\!2JD_{rr}}$, as well as the typical total circulation ${\Gamma\!\sim\!2\pi F(J_{0},t)J_{0}}$. We obtain
$D_{rr}^{2}\sim\frac{J_{0}}{32\pi}\,\gamma\,F(J_{0},t)\sim\frac{1}{64\pi^{2}}\,\Gamma\gamma\,\sim\frac{1}{64\pi^{2}}\,N\,\gamma^{2}.$ (82)
Of course, following our numerous heuristic approximations, the prefactor of Eq. (82) is highly uncertain. Nonetheless, this should be compared with the Taylor–McNamara formula ${D_{rr}^{2}\!=\!N\gamma^{2}/(64\pi^{3})}$ Taylor and McNamara (1971). More generally, Eq. (79) yields a self-consistent equation for the diffusion coefficient which connects the Landau form (${k\Delta\Omega T_{k}\!\gg\!1}$) and the Taylor–McNamara form (${k\Delta\Omega T_{k}\!\ll\!1}$).

### v2

@@ I.1 Regularised Landau equation from individual stochastic equations
$\quad\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\left\langle\mathrm{e}^{\mathrm{i}k\sqrt{D_{\star}/(4J_{\star}^{2})}\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}[\eta(t-t^{\prime})-\eta_{1}(t-t^{\prime})]}\right\rangle$
$\quad\times\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)F(J_{1},t)\bigg],$ (99)
where we introduced the frequency difference, $\Delta\Omega(J,J_{1})\!=\!\Omega(J)\!-\!\Omega(J_{1})$}. Now, if we assume a Gaussian distribution for the fluctuations, we have Dubin (2003)
$\left\langle\mathrm{e}^{\mathrm{i}k\sqrt{D_{\star}/(4J_{\star}^{2})}\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}\eta(t-t^{\prime})}\right\rangle=\mathrm{e}^{-k^{2}D_{\star}/(8J_{\star}^{2})\tau},$ (100)
@@
$\frac{\partial F(J,t)}{\partial t}=4\pi\gamma\frac{\partial}{\partial J}\bigg[\!\int_{0}^{+\infty}\!\!\!\!\!\!\!\!\!\mathrm{d}J_{1}\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}|U_{k}(J,J_{1})|^{2}$ (101)
$\quad\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\mathrm{e}^{-\tau/T_{k}}\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)F(J_{1},t)\bigg],$
where we introduced the (scale-dependent) decorrelation time
$T_{k}\!=\!\frac{4J_{\star}^{2}}{D_{\star}k^{2}}.$ (102)
Integrating Eq. (101) over $\tau$, we obtain
$\frac{\partial F(J,t)}{\partial t}=4\pi^{2}\gamma\frac{\partial}{\partial J}\bigg[\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (103)
$\times\delta_{T_{k}}\big(k\,\Delta\Omega[J,J_{1}]\big)\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)F(J_{1},t)\bigg],$
with $\delta_{T_{k}}(\omega)$} the broadened Dirac delta from Eq. (31). [^10] In the main text, we implicitly assumed $T_{k}\!=\!T_{\mathrm{reg}}/k$}. This leads to the identity $\delta_{T_{k}}(k\,\omega)\!=\!\delta_{T_{\mathrm{reg}}}(\omega)/|k|$}, similar to the one satisfied by the usual Dirac delta. From this identity results the broadened Landau equation (Eq. 32). Here, the function $\delta_{T_{k}}(k\,\omega)$} has a different dependence with $k$, so the broadened Landau equation (Eq. 103) is slightly different. See Chavanis and Lemou (2007) for another regularisation of the Dirac delta obtained by integrating over a finite period of time. Explicitly, the kinetic equation reads
$\frac{\partial F(J,t)}{\partial t}=4\pi\gamma\frac{\partial}{\partial J}\bigg[\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (104)
$\times\frac{T_{k}}{1\!+\!(k\,\Delta\Omega[J,J_{1}]\,T_{k})^{2}}\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)F(J_{1},t)\bigg].$
If we let $T_{k}\!\rightarrow\!+\infty$}, or if $k\,\Delta\Omega\,T_{k}\!\gg\!1$}, which both amount to neglecting the fluctuations in particles individual trajectories, Eq. (104) becomes the Landau equation (Eq. 14). This is valid if we are far from the extremum of $\Omega(J)$}.
However, when $J$ is close to $J_{\star}$, we are faced with the opposite limit, namely $k\,\Delta\Omega\,T_{k}\!\ll\!1$}. In that case, the kinetic equation (Eq. 104) reduces to
$\frac{\partial F(J,t)}{\partial t}={}$ $4\pi\gamma\frac{\partial}{\partial J}\bigg[\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
$\times T_{k}\,\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)F(J_{1},t)\bigg].$ (105)
I.2 Diffusion coefficient at the frequency extremum
@@
$D(J)=8\pi^{2}\gamma\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
$\times\,\delta_{T_{k}}\big(k\,\Delta\Omega[J,J_{1}]\big)F(J_{1},t).$ (106)
@@
As expected, when $T_{k}\!\rightarrow\!+\infty$}, this equation reduces to the usual Landau diffusion coefficient (Eq. 17), which scales like 1/N{1/N} via the prefactor $\gamma$. This limit is relevant if we are sufficiently far from the extremum of frequency. In that case Chavanis and Lemou (2007); Chavanis (2023), we get
$D(J)=8\pi^{2}\gamma\!\sum_{\mathrm{r}}\!\sum_{k>0}k\,|U_{k}(J,J_{\mathrm{r}})|^{2}\frac{F(J_{\mathrm{r}},t)}{|\Omega^{\prime}(J_{\mathrm{r}})|},$ (107)
where $J_{\mathrm{r}}$ are the actions of the resonant particles such that $\Omega(J_{\mathrm{r}})\!=\!\Omega(J)$} (see Eq. 23).
When $J\!\to\!J_{\star}$}, following the same argument as in Eq. (24b), the resonance condition has two roots. Assuming that the frequency profile is quadratic about $J_{\star}$, these two roots are $J_{\mathrm{r}}\!=\!J$} and $J_{\mathrm{r}}\!=\!2J_{\star}\!-\!J$}, which are symmetric with respect to $J_{\star}$. As such, when $J\!\to\!J_{\star}$}, we obtain
$D(J)\simeq 16\pi^{2}\gamma\sum_{k>0}k\,|U_{k}(J_{\star},J_{\star})|^{2}\frac{F(J_{\star},t)}{|\Omega^{\prime}(J)|}.$ (108)
We see that the diffusion coefficient diverges at the extremum of frequency, because of the factor $1/|\Omega^{\prime}(J)|$}. This is the same result as the one obtained in Eq. (25b). When accounting for resonance broadening, this divergence is regularised by the finite value of $T_{k}$, just like it was regularised by $T_{\mathrm{reg}}$ in Eq. (33), as visible in Fig. 6.111111Similarly, in the absence of softening, $\sum_{k}k|U_{k}(J_{\star},J_{\star})|^{2}$} as present in Eq. (108), diverges logarithmically on small scales (see Fig. 4). Fortunately, given that $T_{k}$ also depends on $k$ (see Eq. 102), one can show that resonance broadening also regularises this small scale divergence, see Appendix H.1. Unfortunately, the regularisation from Eq. (32) in the main text still involves the adjustable parameter, $T_{\mathrm{reg}}$, that remains to be determined. In Appendix H, we have estimated $T_{\mathrm{reg}}$ by heuristically relating $T_{\mathrm{reg}}$ to $D_{\star}$ through Eq. (90). However, this equation was not given a precise justification.
In the present approach, $T_{k}$ is related to $D_{\star}$ through the similar Eq. (102), which now arises from a simple stochastic model. Like in Appendix H, one could then hope to obtain the value of $T_{k}$ self-consistently through $D_{\star}$ (see Eq. 102), by applying Eq. (106) at $J\!=\!J_{\star}$}. Indeed, this leads to a self-consistent relation of the form
$D_{\star}=f[D_{\star}],$ (109)
with the function
$f[D_{\star}]=8\pi^{2}\gamma{}$ $\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}J_{1}\,\sum_{k>0}k^{2}|U_{k}(J_{\star},J_{1})|^{2}$
$\times\delta_{T_{k}}\big(k\,\Delta\Omega[J_{\star},J_{1}]\big)F(J_{1},t),$ (110)
where $T_{k}\!=\!T_{k}[D_{\star}]$} follows from Eq. (102). In the present heuristic approach, the self-consistent solution of Eq. (109) should determine $D_{\star}$, hence quantifying the level of resonance broadening occurring in the system.
In Fig. 19, we illustrate the function $D_{\star}\!\mapsto\!f[D_{\star}]$}, with the associated fixed-point search. [^12] This function has the same asymptotic behaviours, $f(D_{\star})\!\propto\!D_{\star}^{-1/2}$} for $D_{\star}\!\ll\!D_{\star}^{\mathrm{t}}$} and $f(D_{\star})\!\propto\!D_{\star}^{-1}$} for $D_{\star}\!\gg\!D_{\star}^{\mathrm{t}}$} with $D_{\star}^{\mathrm{t}}\!\sim\!J_{\star}^{2}/T_{\mathrm{reg}}^{\mathrm{c}}\!\sim\!J_{\star}^{4}|\Omega_{\star}^{\prime\prime}|$}, as the equivalent function studied in Appendix H with a simpler model. Therefore, the discussion of Appendix H on the different regimes also applies to the present situation.
Figure 19: Illustration of the self-consistent relation, $D_{\star}\!=\!f(D_{\star})$} from Eq. (110), for profile (1), here using the rescaled diffusion coefficient, $\mathcal{D}_{\star}$, following Eq. (20b). Here, $\mathcal{D}_{\star}^{\mathrm{t}}$ illustrates the transition between the two regimes investigated in Appendix H.2. For the case N=2 000{N\!=\!2\,000} and q=0.0001{q\!=\!0.0001}, the fixed-point solution is found for $\mathcal{D}_{\mathrm{sol}}\!\simeq\!2.6\!\times\!10^{6}$}. This is $\sim\!$} 20x too large compared to the $N$-body value estimated in Fig. 5.
In that figure, considering the same case as in Fig. 5, we find the fixed-point solution, $\mathcal{D}_{\mathrm{sol}}\!\simeq\!2.6\!\times\!10^{6}$}. This is $\sim\!$} 20x larger than the value for $\mathcal{D}_{\star}$ effectively measured in the numerical simulations. This shows the limitation of the simple stochastic model presented in this Appendix. Surely a refined description of the present resonant broadening is required, e.g., by considering more intricate individual stochastic equations of motion (Eq. 97).
I.3 Link with previous works
The simplified model from Eq. (103) allows us to draw some connections between the Landau diffusion coefficient of sheared flows Chavanis (1998, 2001); Dubin (2003) and the Taylor–McNamara diffusion coefficient of unsheared flows Taylor and McNamara (1971). These correspond to two important limits considered in the literature, each with different scalings with respect to $N$.
For the sake of simplicity, let us assume that the vorticity and frequency profiles are approximately flat on a domain of size $J_{\star}$, writing $F\!\simeq\!F_{\star}$} and assuming $k\,\Delta\Omega\,T_{k}\!\ll\!1$}. From Eq. (106), we obtain
$D^{2}\simeq 32\pi\gamma\,J_{\star}^{2}F_{\star}\!\!\int_{0}^{J_{\star}}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}|U_{k}(J,J_{1})|^{2}.$ (111)
@@ I.3 Link with previous works
Since $\gamma\!\sim\!1/N$}, we recover the Taylor–McNamara scaling Taylor and McNamara (1971), $D\!\sim\!1/\sqrt{N}$}. If we account for an active fraction of particles, as in the main text, this scaling becomes $D\!\sim\!q/\sqrt{N}$}.
@@
The integral in Eq. (111) can be estimated analytically when using the (unsoftened) interaction potential from Eq. (52). It leads to
$D^{2}\simeq\frac{2J_{\star}^{3}}{\pi}\,\gamma\,F_{\star},$ (112)
@@
up to logarithmic corrections. To make the connection with the results of Taylor–McNamara Taylor and McNamara (1971), we introduce the original isotropic diffusion coefficient in the physical space, (x,y){(x,y)}, via $D_{\mathrm{iso}}\!\sim\!D/(2J_{\star})$}, as well as the total circulation $\Gamma\!\sim\!2\pi F_{\star}J_{\star}$}. We obtain
$D_{\mathrm{iso}}^{2}\simeq\frac{J_{\star}}{2\pi}\,\gamma\,F_{\star}\sim\frac{1}{4\pi^{2}}\,\Gamma\gamma\,\sim\frac{1}{4\pi^{2}}\,N\,\gamma^{2}.$ (113)
Of course, following our numerous approximations, the prefactor of Eq. (113) is uncertain by a factor of order unity. Nonetheless, this should be compared with the Taylor–McNamara formula, $D_{\mathrm{TN}}^{2}\!=\!N\gamma^{2}/(16\pi^{3})$} Taylor and McNamara (1971). Our estimate of $D_{\mathrm{iso}}$ turns out to be larger than $D_{\mathrm{TN}}$ by a factor $\sqrt{4\pi}\!\sim\!3.5$}.